OpenStudy (anonymous):

Factor : 2a^2 - 5a + 4

OpenStudy (anonymous):

-3a^2 + 4 ? i think thats it

OpenStudy (anonymous):

its a quadratic in a so jst solve it

OpenStudy (anonymous):

2a^2 - 5a + 4 = 2 (a^2 - 5a/2) + 4 = 2 (a^2 - 5a/2 + 25/16) + 4 - 25/4 = 2 (a - 5/4)^2 - (5/2)^2 = (2a - 5/2 - (5/2)^0.5)(2a - 5/2 + (5/2)^0.5)

OpenStudy (anonymous):